10. Wanda Lux must select a supplier for a plastic bottle and proprietary dispenser for its new hair shampoo. Three suppliers have placed bids; at Wandas request, all bids are for a shipping quantity (Q) of 20,000 bottles with annual requirements (D) of 40,000 units. Wandas factory operates 250 days a year. The following table shows each suppliers price, estimated annual freight costs, and current lead times; management has added estimates for holding costs and administrative oversight costs for each supplier Supplier Dover Plastics Evan & Sons Farley, Inc. Freight Price/Unit Annual Holding Lead Time Annual Administrative Costs $4,500 $3,500 $3,000 $5.15 $5.05 $5.00 Cost/Unit (H $1.03 $1.01 $1.00 (L) (days) 20 15 Costs $6,000 $3,000 $5,000 a. Which of the three suppliers will provide the lowest annual cost to Wanda Lux? Using (1) your response rounded to the nearest whole number.) is the lowest cost alternative, with annual total costs to Wanda Lux of S (Enter Beyond costs, however, Wanda has three other criteria considered important in the selection of a supplier. The following table shows all the criteria, their weights, and the scores for all of them except total costs, where a score of 1.0 indicates poor and 10 indicates superior. Because all three suppliers have done business with Wanda Lux before, management will assign a score of 10 to the supplier with the lowest total annual cost, a score of 8.5 for the next lowest cost, and a score of 7.0 for the worst cost of the three Rating Evan Criteria 1. Total Cost 2. Consistent Quality 3. On-Time Delivery 4. Environment Wei Dover Fari 30 20 20 b. Given Wandas criteria and weighting system, which supplier should Wanda award the contract to? whose total weighted score is (Enter your Wanda should award the contract to (2) response as an integer.) 1) Farley (2) O Farley O Dover Evan O Evan
Answer #1

a) Number of shipments per year = D/Q = 40000/20000 = 2

Total annual cost of Dover Plastics = Freight costs + Item cost + Annual holding cost of cycle stock + Annual holding cost of transit stock + Annual administrative costs

= 4500 + 5.15*40000 + (20000/2)*1.03 + 40000*1.03*20/250 + 6000

= $ 230,096

Total annual cost of Evan & Sons = Freight costs + Item cost + Annual holding cost of cycle stock + Annual holding cost of transit stock + Annual administrative costs

= 3500 + 5.05*40000 + (20000/2)*1.01 + 40000*1.01*15/250 + 3000

= $ 221,024

Total annual cost of Farley, Inc = Freight costs + Item cost + Annual holding cost of cycle stock + Annual holding cost of transit stock + Annual administrative costs

= 3000 + 5.00*40000 + (20000/2)*1.00 + 40000*1.00*17/250 + 5000

= $ 220,720

Using (1) Farley, Inc. is the lowest cost alternative, with annual total costs to Wanda Lux of $ 220,720

b) Critieria table is following:

Rating
Criteria Weight Dover Evan Farley
1. Total Cost 30 7 8.5 10
2. Consistent Quality 30 7 9 8
3. Ontime delivery 20 7 8 7
4. Environment 20 9 9 8
Total Weighted Score 7.4 8.65 8.4

Total weighted score is calculated as =SUMPRODUCT(Weight*Rating)/SUM(Weights)

Total weighted score of Dover = (30*7+30*7+20*7+20*9)/(30+30+20+20) = 7.40

Total weighted score of Evan = (30*8.5+30*9+20*8+20*9)/(30+30+20+20) = 8.65

Total weighted score of Farley = (30*10+30*8+20*7+20*8)/(30+30+20+20) = 8.40

Wanda should award the contract to (2) Evan , whose total weighted score is 8.65